A. Calculate the pH of a buffer made from 0.28 M HNO2 and 0.23




M NO2-. The ka for HNO2 is 4.6 x 10-4.





1
Expert's answer
2022-04-29T14:54:04-0400

Solution:

The Henderson-Hasselbalch equation is commonly used to calculate the pH of a buffer solution from the concentration of the buffer components:



where pKa is the acid dissociation constant and [base] and [acid] are the base and acid concentrations, respectively, in the chemical equation.

In this specific reaction, the base is NO2, and the acid is HNO2.

Therefore,

pH = −log(4.6×10−4) + log(0.23 / 0.28) = 3.3372 + (−0.0854) = 3.2518 = 3.25

pH = 3.25


Answer: The pH of a buffer is 3.25